Consider a day for a local junk removal company. The inputs the company uses to haul junk are two employees each paid an hourly wage of $15 per hour and a moving truck, which costs $100 (splitting the cost of ownership over periods). Each hauling job takes 1 hour.

  1. Are labor hours a fixed or variable input? Explain.
  2. Is the moving truck a fixed or variable input? Explain.
  3. Determine the total cost (TC), total fixed cost (TFC), total variable cost (TVC), average total cost (ATC), average fixed cost (AFC), average variable cost (AVC), and the marginal cost (MC) of trimming bushes for residential lawns in the table below.

  1. Graph the ATC, AFC, AVC, and MC curves.

Jobs

TC

TFC

TVC

ATC

AFC

AVC

MC

0

1

2

3

4

5

6

7

8

9

10
Answer #1

a) labur hours are variable inputs as it depends on the number of hours worked and it increase with increase in number of hours worked.

b)moving truck is a fixed input as it is independent of number of hours worked.

c)

Jobs TC TFC TVC = (30*hours) ATC AFC AVC MC = ΔTC / ΔQ
0 100 100 0
1 130 100 30 130.00 100.00 30 30
2 160 100 60 80.00 50.00 30 30
3 190 100 90 63.33 33.33 30 30
4 220 100 120 55.00 25.00 30 30
5 250 100 150 50.00 20.00 30 30
6 280 100 180 46.67 16.67 30 30
7 310 100 210 44.29 14.29 30 30
8 340 100 240 42.50 12.50 30 30
9 370 100 270 41.11 11.11 30 30
10 400 100 300 40.00 10.00 30 30

d) 140 120 100 80 ATC MCAVCATCAFC 60 AFC AVC 40 * MC = ATC / 40 20 0 0 2 4 6 8 10 12 Jobs Hours